# Stability of Poiseuille Flow of Navier-Stokes Equations on $\mathbb{R}^{2}$

### v1

@@ 3.7 Bound on $T_{6}$
$T_{6,HL,L,LH}^{x}$ $=-2c_{\gamma}\iint 1_{|k-k^{\prime}|>\frac{|k|}{2}}1_{|k|<\nu^{-\frac{1}{3}}}1_{|k|\leq|k^{\prime}|}\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}\gamma_{k}\mathrm{Re}\langle\partial_{y}\psi_{k}\overline{\psi_{k-k^{\prime}}},i(k-k^{\prime})\omega_{k^{\prime}}\rangle\mathrm{d}k\mathrm{d}k^{\prime}$
$-2c_{\gamma}\iint 1_{|k-k^{\prime}|>\frac{|k|}{2}}1_{|k|<\nu^{-\frac{1}{3}}}1_{|k|\leq|k^{\prime}|}\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}\gamma_{k}\mathrm{Re}\langle\psi_{k}\partial_{y}\overline{\psi_{k-k^{\prime}}},i(k-k^{\prime})\omega_{k^{\prime}}\rangle\mathrm{d}k\mathrm{d}k^{\prime}$
@@
We treat $T_{6,HL,L,LH}^{x,1}$ first. Applying Cauchy-Schwarz inequality in $y$, Young’s convolution inequality and Lemma 3.4, we obatain
$|T_{6,HL,L,LH}^{x,1}|$ $\lesssim\iint\langle c\lambda_{k-k^{\prime}}t\rangle^{J}\langle k-k^{\prime}\rangle^{m}\|(k-k^{\prime})\psi_{k-k^{\prime}}\|_{2}\langle c\lambda_{k^{\prime}}t\rangle^{J}\langle k^{\prime}\rangle^{m}\|\omega_{k^{\prime}}\|_{2}\nu^{-\frac{2}{3}}\|\partial_{y}\psi_{k}\|_{\infty}\mathrm{d}k\mathrm{d}k^{\prime}$
$\lesssim\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\|\nabla_{k}\psi_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\|\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k$
$\lesssim\nu^{-\frac{7}{6}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}.$
For $T_{6,HL,L,LH}^{x,2}$, by Cauchy-Schwarz inequality in $y$, Young’s convolution inequality, and $\|k\partial_{y}\psi_{k}\|_{2}\lesssim\|\Delta_{k}\psi_{k}\|_{2}$, we have
$|T_{6,HL,L,LH}^{x,2}|$ $\lesssim\iint 1_{|k|<\nu^{-\frac{1}{3}}}\langle c\lambda_{k-k^{\prime}}t\rangle^{J}\langle k-k^{\prime}\rangle^{m}\|(k-k^{\prime})\partial_{y}\psi_{k-k^{\prime}}\|_{2}\langle c\lambda_{k^{\prime}}t\rangle^{J}\langle k^{\prime}\rangle^{m}\|\omega_{k^{\prime}}\|_{2}\nu^{-\frac{2}{3}}\|\psi_{k}\|_{\infty}\mathrm{d}k\mathrm{d}k^{\prime}$
$\lesssim\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\|\Delta_{k}\psi_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\|\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int_{|k|<\nu^{-\frac{1}{3}}}\nu^{-\frac{2}{3}}\|\psi_{k}\|_{\infty}\mathrm{d}k.$
To estimate the last term, we use Gagliardo-Nirenberg-Sobolev inequality and Hölder’s inequality,
$\int_{|k|<\nu^{-\frac{1}{3}}}\nu^{-\frac{2}{3}}\|\psi_{k}\|_{\infty}\mathrm{d}k\lesssim\int_{|k|<\nu^{-\frac{1}{3}}}\nu^{-\frac{2}{3}}|k|^{-\frac{1}{2}}|k|^{\frac{1}{2}}\|\psi_{k}\|_{2}^{\frac{1}{2}}\|\partial_{y}\psi_{k}\|_{2}^{\frac{1}{2}}\mathrm{d}k\lesssim\int_{|k|<\nu^{-\frac{1}{3}}}\nu^{-\frac{2}{3}}|k|^{-\frac{1}{2}}\mathrm{d}k\sup_{k}\|\nabla_{k}\psi_{k}\|_{2}\lesssim\nu^{-\frac{5}{6}}\mathcal{E}^{\frac{1}{2}}.$
@@
$|T_{6,HL,L,LH}^{x,2}|\lesssim\nu^{-\frac{1}{3}}\mathcal{D}\nu^{-\frac{5}{6}}\mathcal{E}^{\frac{1}{2}}=\nu^{-\frac{7}{6}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}.$
@@
$|T_{6,HL,L,LH}^{y}|\lesssim\iint 1_{|k|<\nu^{-\frac{1}{3}}}\langle c\lambda_{k-k^{\prime}}t\rangle^{J}\langle k-k^{\prime}\rangle^{m}\|(k-k^{\prime})\partial_{y}\psi_{k-k^{\prime}}\|_{2}\langle c\lambda_{k^{\prime}}t\rangle^{J}\langle k^{\prime}\rangle^{m}\|\omega_{k^{\prime}}\|_{2}\nu^{-\frac{2}{3}}\|\psi_{k}\|_{\infty}\mathrm{d}k\mathrm{d}k^{\prime}\lesssim\nu^{-\frac{7}{6}}\mathcal{D}\mathcal{E}^{\frac{1}{2}},$
where we have used $|k^{\prime}|\lesssim|k-k^{\prime}|$. Now we treat the last term $T_{6,HL,L,HL}$. As before, we separate it into two parts, $T_{6,HL,L,HL}\eqqcolon T_{6,HL,L,HL}^{x}+T_{6,HL,L,HL}^{y}$. We still separate $T_{6,HL,L,HL}^{x}$ into two parts by integration by parts,
$T_{6,HL,L,HL}^{x}$ $=-2c_{\gamma}\iint 1_{|k-k^{\prime}|>\frac{|k|}{2}}1_{|k|<\nu^{-\frac{1}{3}}}1_{|k|>|k^{\prime}|}\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}\gamma_{k}\mathrm{Re}\langle\partial_{y}\psi_{k}\overline{\psi_{k-k^{\prime}}},i(k-k^{\prime})\omega_{k^{\prime}}\rangle\mathrm{d}k\mathrm{d}k^{\prime}$
$-2c_{\gamma}\iint 1_{|k-k^{\prime}|>\frac{|k|}{2}}1_{|k|<\nu^{-\frac{1}{3}}}1_{|k|>|k^{\prime}|}\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}\gamma_{k}\mathrm{Re}\langle\psi_{k}\partial_{y}\overline{\psi_{k-k^{\prime}}},i(k-k^{\prime})\omega_{k^{\prime}}\rangle\mathrm{d}k\mathrm{d}k^{\prime}$
@@
To estimate $T_{6,HL,L,HL}^{x,1}$, notice $|k^{\prime}|\lesssim|k-k^{\prime}|$ on the domain of integration, and we apply Cauchy-Schwarz inequality in $y$, Young’s convolution inequality,
@@
$|T_{6,HL,L,HL}^{x,1}|$ $\lesssim\iint 1_{|k^{\prime}|\lesssim\nu^{-\frac{1}{3}}}\langle c\lambda_{k}t\rangle^{J}\langle k\rangle^{m}\nu^{-\frac{1}{3}}\gamma_{k}^{\frac{1}{2}}\|\partial_{y}\psi_{k}\|_{2}\langle c\lambda_{k-k^{\prime}}t\rangle^{J}\langle k-k^{\prime}\rangle^{m}|k-k^{\prime}|^{\frac{1}{2}}\|(k-k^{\prime})\psi_{k-k^{\prime}}\|_{\infty}|k^{\prime}|^{-\frac{1}{2}}\|\omega_{k^{\prime}}\|_{2}\mathrm{d}k\mathrm{d}k^{\prime}$
@@
$\lesssim\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\nu^{-\frac{2}{3}}\gamma_{k}\|\nabla_{k}\psi_{k}\|_{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}|k|\|\nabla_{k}\psi_{k}\|_{\infty}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int_{|k|\lesssim\nu^{-\frac{1}{3}}}|k|^{-\frac{1}{2}}\|\omega_{k}\|_{2}\mathrm{d}k.$
For the last term, by Hölder’s inequality,
@@
$\int_{|k|\lesssim\nu^{-\frac{1}{3}}}|k|^{-\frac{1}{2}}\|\omega_{k}\|_{2}\mathrm{d}k\lesssim\int_{|k|\lesssim\nu^{-\frac{1}{3}}}|k|^{-\frac{1}{2}}\mathrm{d}k\sup_{k}\|\omega_{k}\|_{2}\lesssim\nu^{-\frac{1}{6}}(\nu^{-1}\mathcal{D})^{\frac{1}{2}}=\nu^{-\frac{2}{3}}\mathcal{D}^{\frac{1}{2}}.$

### v4

@@ 3.5 Bound on $T_{6}$
$T_{6,HL,L,LH}^{x}=$ $-2c_{\gamma}\int_{0}^{t}\iint_{|k-k^{\prime}|>\frac{|k|}{2},|k|<\nu^{-\frac{1}{3}},|k|\leq|k^{\prime}|}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}\gamma_{k}\mathrm{Re}\langle\partial_{y}\psi_{k}\overline{\psi_{k-k^{\prime}}},i(k-k^{\prime})\omega_{k^{\prime}}\rangle\mathrm{d}k\mathrm{d}k^{\prime}\mathrm{d}s$
$-2c_{\gamma}\int_{0}^{t}\iint_{|k-k^{\prime}|>\frac{|k|}{2},|k|<\nu^{-\frac{1}{3}},|k|\leq|k^{\prime}|}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}\gamma_{k}\mathrm{Re}\langle\psi_{k}\partial_{y}\overline{\psi_{k-k^{\prime}}},i(k-k^{\prime})\omega_{k^{\prime}}\rangle\mathrm{d}k\mathrm{d}k^{\prime}\mathrm{d}s$
@@
We treat $T_{6,HL,L,LH}^{x,1}$ first. Applying Cauchy-Schwarz inequality in $y$, Young’s convolution inequality, Hölder’s inequality in $s$ and Lemma 3.5, we obatain
$|T_{6,HL,L,LH}^{x,1}|$ $\lesssim\int_{0}^{t}\iint_{\mathbb{R}^{2}}\langle c\lambda_{k-k^{\prime}}s\rangle^{J}\langle k-k^{\prime}\rangle^{m}\gamma_{k-k^{\prime}}^{\frac{1}{2}}\|(k-k^{\prime})\psi_{k-k^{\prime}}\|_{2}\langle c\lambda_{k^{\prime}}s\rangle^{J}\langle k^{\prime}\rangle^{m}\gamma_{k^{\prime}}^{\frac{1}{2}}\|\omega_{k^{\prime}}\|_{2}\|\partial_{y}\psi_{k}\|_{\infty}\mathrm{d}k\mathrm{d}k^{\prime}\mathrm{d}s$
$\lesssim\int_{0}^{t}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}\|\nabla_{k}\psi_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}\|\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int_{\mathbb{R}}\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k\mathrm{d}s$
$\lesssim\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}}\left(\int_{0}^{t}\nu^{-1}\widetilde{\mathcal{D}}\mathrm{d}s\right)^{\frac{1}{2}}\left(\int_{0}^{t}\left(\int_{\mathbb{R}}\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k\right)^{2}\mathrm{d}s\right)^{\frac{1}{2}}$
$\lesssim\nu^{-\frac{7}{6}}\mathcal{D}(t)\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}}.$
Turning to $T_{6,HL,L,LH}^{x,2}$, we use Cauchy-Schwarz inequality in $y$, Young’s convolution inequality, together with $\|k\partial_{y}\psi_{k}\|_{2}\lesssim\|\Delta_{k}\psi_{k}\|_{2}=\|\omega_{k}\|_{2}$,
$|T_{6,HL,L,LH}^{x,2}|$ $\lesssim\int_{0}^{t}\iint_{|k|<\nu^{-\frac{1}{3}}}\langle c\lambda_{k-k^{\prime}}s\rangle^{J}\langle k-k^{\prime}\rangle^{m}\gamma_{k-k^{\prime}}^{\frac{1}{2}}\|(k-k^{\prime})\partial_{y}\psi_{k-k^{\prime}}\|_{2}\langle c\lambda_{k^{\prime}}s\rangle^{J}\langle k^{\prime}\rangle^{m}\gamma_{k^{\prime}}^{\frac{1}{2}}\|\omega_{k^{\prime}}\|_{2}\|\psi_{k}\|_{\infty}\mathrm{d}k\mathrm{d}k^{\prime}\mathrm{d}s$
$\lesssim\int_{0}^{t}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}\|\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}\|\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int_{|k|<\nu^{-\frac{1}{3}}}\|\psi_{k}\|_{\infty}\mathrm{d}k\mathrm{d}s.$
To estimate the last term, we use Gagliardo-Nirenberg-Sobolev inequality,
$\int_{|k|<\nu^{-\frac{1}{3}}}\|\psi_{k}\|_{\infty}\mathrm{d}k\lesssim\int_{|k|<\nu^{-\frac{1}{3}}}|k|^{-\frac{1}{2}}|k|^{\frac{1}{2}}\|\psi_{k}\|_{2}^{\frac{1}{2}}\|\partial_{y}\psi_{k}\|_{2}^{\frac{1}{2}}\mathrm{d}k\lesssim\int_{|k|<\nu^{-\frac{1}{3}}}|k|^{-\frac{1}{2}}\mathrm{d}k\cdot\sup_{k\in\mathbb{R}}\|\nabla_{k}\psi_{k}\|_{2}\lesssim\nu^{-\frac{1}{6}}\mathcal{E}^{\frac{1}{2}}.$
@@
$|T_{6,HL,L,LH}^{x,2}|\lesssim\int_{0}^{t}\nu^{-1}\widetilde{\mathcal{D}}\nu^{-\frac{1}{6}}\mathcal{E}^{\frac{1}{2}}\mathrm{d}s\lesssim\nu^{-\frac{7}{6}}\mathcal{D}(t)\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}}.$
@@
$|T_{6,HL,L,LH}^{y}|$ $\lesssim\int_{0}^{t}\iint_{|k|<\nu^{-\frac{1}{3}}}\langle c\lambda_{k-k^{\prime}}s\rangle^{J}\langle k-k^{\prime}\rangle^{m}\gamma_{k-k^{\prime}}^{\frac{1}{2}}\|(k-k^{\prime})\partial_{y}\psi_{k-k^{\prime}}\|_{2}\langle c\lambda_{k^{\prime}}s\rangle^{J}\langle k^{\prime}\rangle^{m}\gamma_{k^{\prime}}^{\frac{1}{2}}\|\omega_{k^{\prime}}\|_{2}\nu^{-\frac{2}{3}}\|\psi_{k}\|_{\infty}\mathrm{d}k\mathrm{d}k^{\prime}\mathrm{d}s$
$\lesssim\nu^{-\frac{7}{6}}\mathcal{D}(t)\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}},$
where we have used $|k^{\prime}|\lesssim|k-k^{\prime}|$ since $|k-k^{\prime}|>\frac{|k|}{2}$. Now we treat the last term $T_{6,HL,L,HL}$. As before, we separate it into two parts, $T_{6,HL,L,HL}\eqqcolon T_{6,HL,L,HL}^{x}+T_{6,HL,L,HL}^{y}$. We still separate $T_{6,HL,L,HL}^{x}$ into two parts by integration by parts,
$T_{6,HL,L,HL}^{x}=$ $-2c_{\gamma}\int_{0}^{t}\iint_{|k-k^{\prime}|>\frac{|k|}{2},|k|<\nu^{-\frac{1}{3}},|k|>|k^{\prime}|}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}\gamma_{k}\mathrm{Re}\langle\partial_{y}\psi_{k}\overline{\psi_{k-k^{\prime}}},i(k-k^{\prime})\omega_{k^{\prime}}\rangle\mathrm{d}k\mathrm{d}k^{\prime}\mathrm{d}s$
$-2c_{\gamma}\int_{0}^{t}\iint_{|k-k^{\prime}|>\frac{|k|}{2},|k|<\nu^{-\frac{1}{3}},|k|>|k^{\prime}|}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}\gamma_{k}\mathrm{Re}\langle\psi_{k}\partial_{y}\overline{\psi_{k-k^{\prime}}},i(k-k^{\prime})\omega_{k^{\prime}}\rangle\mathrm{d}k\mathrm{d}k^{\prime}\mathrm{d}s$
@@
To estimate $T_{6,HL,L,HL}^{x,1}$, notice $|k^{\prime}|\lesssim|k-k^{\prime}|$ on the domain of integration, and we apply Cauchy-Schwarz inequality in $y$, Young’s convolution inequality, Hölder’s inequality in $s$ and Lemma 3.6,
@@
$|T_{6,HL,L,HL}^{x,1}|\!$ $\lesssim\!\int_{0}^{t}\!\iint_{|k^{\prime}|\lesssim\nu^{-\frac{1}{3}}}\!\langle c\lambda_{k}s\rangle^{J}\langle k\rangle^{m}\nu^{-\frac{1}{3}}\gamma_{k}^{\frac{1}{2}}\|\partial_{y}\psi_{k}\|_{2}\langle c\lambda_{k-k^{\prime}}s\rangle^{J}\langle k-k^{\prime}\rangle^{m}|k-k^{\prime}|^{\frac{1}{2}}\|(k-k^{\prime})\psi_{k-k^{\prime}}\|_{\infty}|k^{\prime}|^{-\frac{1}{2}}\|\omega_{k^{\prime}}\|_{2}\mathrm{d}k\mathrm{d}k^{\prime}\mathrm{d}s$
@@
$\lesssim\int_{0}^{t}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\nu^{-\frac{2}{3}}\gamma_{k}\|\nabla_{k}\psi_{k}\|_{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}|k|\|\nabla_{k}\psi_{k}\|_{\infty}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int_{|k|\lesssim\nu^{-\frac{1}{3}}}|k|^{-\frac{1}{2}}\|\omega_{k}\|_{2}\mathrm{d}k\mathrm{d}s$
@@
$\lesssim\sup_{s\in[0,t]}(\nu^{-\frac{2}{3}}\mathcal{E}(s))^{\frac{1}{2}}\left(\int_{0}^{t}\nu^{-\frac{1}{3}}\widetilde{\mathcal{D}}\mathrm{d}s\right)^{\frac{1}{2}}\left(\int_{0}^{t}\left(\int_{|k|\lesssim\nu^{-\frac{1}{3}}}|k|^{-\frac{1}{2}}\|\omega_{k}\|_{2}\mathrm{d}k\right)^{2}\mathrm{d}s\right)^{\frac{1}{2}}.$
For the last term, by Minkowski’s inequality for integrals,
$\left(\int_{0}^{t}\left(\int_{|k|\lesssim\nu^{-\frac{1}{3}}}|k|^{-\frac{1}{2}}\|\omega_{k}\|_{2}\mathrm{d}k\right)^{2}\mathrm{d}s\right)^{\frac{1}{2}}$ $\lesssim\int_{|k|\lesssim\nu^{-\frac{1}{3}}}\left(\int_{0}^{t}\left(|k|^{-\frac{1}{2}}\|\omega_{k}\|_{2}\right)^{2}\mathrm{d}s\right)^{\frac{1}{2}}\mathrm{d}k$
$\lesssim\int_{|k|\lesssim\nu^{-\frac{1}{3}}}\nu^{-\frac{1}{2}}|k|^{-\frac{1}{2}}\mathrm{d}k\cdot\sup_{k\in\mathbb{R}}\left(\int_{0}^{t}\nu\|\omega_{k}\|_{2}^{2}\right)^{\frac{1}{2}}$
$\lesssim\nu^{-\frac{2}{3}}\mathcal{D}^{\frac{1}{2}}.$
Thus
$|T_{6,HL,L,HL}^{x,1}|\lesssim\nu^{-\frac{7}{6}}\mathcal{D}(t)\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}}.$
To handle $T_{6,HL,L,HL}^{x,2}$, still using Cauchy-Schwarz inequality in $y$, Young’s convolution inequality, Hölder’s inequality in $s$, together with $\|k\partial_{y}\psi_{k}\|_{2}\lesssim\|\Delta_{k}\psi_{k}\|_{2}=\|\omega_{k}\|_{2}$,
$|T_{6,HL,L,HL}^{x,2}|$ $\lesssim\int_{0}^{t}\iint_{|k^{\prime}|\lesssim\nu^{-\frac{1}{3}}}\langle c\lambda_{k}s\rangle^{J}\langle k\rangle^{m}\nu^{-\frac{1}{3}}\gamma_{k}^{\frac{1}{2}}|k|^{\frac{1}{2}}\|\psi_{k}\|_{\infty}\langle c\lambda_{k-k^{\prime}}s\rangle^{J}\langle k-k^{\prime}\rangle^{m}\|(k-k^{\prime})\partial_{y}\psi_{k-k^{\prime}}\|_{2}|k^{\prime}|^{-\frac{1}{2}}\|\omega_{k^{\prime}}\|_{2}\mathrm{d}k\mathrm{d}k^{\prime}\mathrm{d}s$
$\lesssim\int_{0}^{t}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\nu^{-\frac{2}{3}}\gamma_{k}|k|\|\psi_{k}\|_{\infty}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\|\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int_{|k|\lesssim\nu^{-\frac{1}{3}}}|k|^{-\frac{1}{2}}\|\omega_{k}\|_{2}\mathrm{d}k\mathrm{d}s$
$\lesssim\sup_{s\in[0,t]}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\nu^{-\frac{2}{3}}\gamma_{k}|k|\|\psi_{k}\|_{\infty}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int_{0}^{t}\nu^{-\frac{1}{3}}\widetilde{\mathcal{D}}\mathrm{d}s\right)^{\frac{1}{2}}\left(\int_{0}^{t}\left(\int_{|k|\lesssim\nu^{-\frac{1}{3}}}|k|^{-\frac{1}{2}}\|\omega_{k}\|_{2}\mathrm{d}k\right)^{2}\mathrm{d}s\right)^{\frac{1}{2}}.$
By Gagliardo-Nirenberg-Sobolev inequality, the first term obeys
$\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\nu^{-\frac{2}{3}}\gamma_{k}|k|\|\psi_{k}\|_{\infty}^{2}\mathrm{d}k\lesssim\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\nu^{-\frac{2}{3}}\gamma_{k}|k|\|\psi_{k}\|_{2}\|\partial_{y}\psi_{k}\|_{2}\mathrm{d}k\lesssim\nu^{-\frac{2}{3}}\mathcal{E}.$
For the third term, it is the same as the third term of $|T_{6,HL,L,HL}^{x,1}|$, so we obtain
$|T_{6,HL,L,HL}^{x,2}|\lesssim\nu^{-\frac{7}{6}}\mathcal{D}(t)\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}}.$
The estimate of $T_{6,HL,L,HL}^{y}$ is similar to $T_{6,HL,L,HL}^{x,2}$, and then we have
$|T_{6,HL,L,HL}^{y}|\lesssim\nu^{-\frac{7}{6}}\mathcal{D}(t)\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}}.$
3.6 Bound on $T_{7}$
Bounding the term is the simplest. We use Cauchy-Schwarz inequality in $y$,
$|T_{7}|\lesssim\sup_{k\in\mathbb{R}}\int_{0}^{t}\int_{\mathbb{R}}\|y\omega_{k}\|_{2}\|\nabla_{k-k^{\prime}}^{\bot}\psi_{k-k^{\prime}}\|_{\infty}\|y\nabla_{k^{\prime}}\omega_{k^{\prime}}\|_{2}\mathrm{d}k^{\prime}\mathrm{d}s.$
By Fubini’s theorem, we are able to integrate with respect to $s$ first, and then apply Hölder’s inequality for the integration over $s$,
$|T_{7}|\lesssim\sup_{k\in\mathbb{R}}\int_{\mathbb{R}}\sup_{s\in[0,t]}\|y\omega_{k}\|_{2}\left(\int_{0}^{t}\|\nabla_{k-k^{\prime}}\psi_{k-k^{\prime}}\|_{\infty}^{2}\mathrm{d}s\right)^{\frac{1}{2}}\left(\int_{0}^{t}\|y\nabla_{k^{\prime}}\omega_{k^{\prime}}\|_{2}^{2}\mathrm{d}s\right)^{\frac{1}{2}}\mathrm{d}k^{\prime}.$
We use Young’s convolution inequality to place the $k$ and $k^{\prime}$ factors into $L^{\infty}$ and the $k-k^{\prime}$ factor into $L^{1}$,
$|T_{7}|\lesssim\sup_{k\in\mathbb{R}}\sup_{s\in[0,t]}\|y\omega_{k}\|_{2}\cdot\int_{\mathbb{R}}\left(\int_{0}^{t}\|\nabla_{k}\psi_{k}\|_{\infty}^{2}\mathrm{d}s\right)^{\frac{1}{2}}\mathrm{d}k\cdot\sup_{k\in\mathbb{R}}\left(\int_{0}^{t}\|y\nabla_{k}\omega_{k}\|_{2}^{2}\mathrm{d}s\right)^{\frac{1}{2}}.$